# Coverage and Bias of Street View Imagery in Mapping the Urban Environment

### v2

@@ Abstract
Street View Imagery (SVI) has emerged as a valuable data form in urban studies, enabling new ways to map and sense urban environments. However, fundamental concerns regarding the representativeness, quality, and reliability of SVI remain underexplored, e.g. to what extent can cities be captured by such data and do data gaps result in bias. This research, positioned at the intersection of spatial data quality and urban analytics, addresses these concerns by proposing a novel and effective method to estimate SVI’s element-level coverage in the urban environment. The method integrates the positional relationships between SVI and target elements, as well as the impact of physical obstructions. Expanding the domain of data quality to SVI, we introduce an indicator system that evaluates the extent of coverage, focusing on the completeness and frequency dimensions. Taking London as a case study, three experiments are conducted to identify potential biases in SVI’s ability to cover and represent urban environmental elements, using building facades as an example. It is found that despite their high availability along urban road networks, Google Street View covers only 62.4% of buildings in the case study area. The average facade coverage per building is 12.4%. SVI tends to over-represent non-residential buildings, thus possibly resulting in biased analyses, and its coverage of environmental elements is position-dependent. The research also highlights the variability of SVI coverage under different data acquisition practices and proposes an optimal sampling interval range of 50-60 m for SVI collection. The findings suggest that while SVI offers valuable insights, it is no panacea – its application in urban research requires careful consideration of data coverage and element-level representativeness to ensure reliable results.
@@ 2.2.1 SVI Data Quality: Availability, Image Quality, and Others
SVI data quality problems can be regarded as an inherent source of uncertainties for SVI in representing urban environment. Among them, the data availability have received more attention, and there are uneven availability distribution of SVI images and services across various geographic scales. Notably, a significant number of cities worldwide still lack SVI services. Cities in Europe and North America enjoy broader SVI coverage, whereas economically underdeveloped regions in Latin America and Africa experience sparse and limited distribution (Bendixen et al., 2023; Hou and Biljecki, 2022; Quinn and Alvarez León, 2019). Moreover, there is spatial heterogeneity of SVI collection within cities where SVI services are available. For example, streets in areas characterized by high traffic volumes, dense populations, or wealthier demographics are more likely to be imaged (Fry et al., 2020). Conversely, smaller towns and rural areas are often overlooked (Szczepańska and Pietrzyk, 2020). Additionally, image availability in informal urban sectors can be compromised by the absence of accessible roads (Chen et al., 2022; He and Li, 2021), which results in the structured missingness problem (Mitra et al., ). In their review papers, Kang et al. (2020) and Biljecki and Ito (2021) point out that SVI is predominantly collected along streets, making it challenging to analyze variations within neighborhood built environments, potentially leading to structural issues in the completeness of information.
@@ 2.2.2 Common Practices in SVI Utilization
Beyond the image sampling methods and intervals, researchers also consider the impact from other parameter selections and practices on SVI utilization. Specifically, these studies have compared different SVI orientations (Kim et al., 2021), different collection positions (lanes and sidewalks) (Ki et al., 2023; Ito et al., 2024), and different image sources and forms (crowdsourced and commercial; perspective and panoramic) (Biljecki et al., 2023b), to see if they affect how SVI summarizes and reflects the same urban environment elements, such as buildings, greenery, and sky. It is reported that though reliability of a single, crowdsourced imagery is comparable to commercial panoramas (Biljecki et al., 2023b), there are significant measurement errors for sidewalks, greenery, and roads between pedestrian and vehicle views (Ki et al., 2023). Additionally, Liu and Sevtsuk (2024) discussed the issues such as lack of clear technical definitions in street attributes extracted from SVI, the lag in CV model performance, and the absence of benchmarks. Although not the focus of this paper, these issues contribute another aspect of uncertainty in the common practices of SVI utilization.
@@ 2.2.3 Complexity of Real-world Environment
Additionally, when mapping buildings or trees in urban environment, obstructions can also be caused by the targeted elements themselves (Yan and Huang, 2022; Raghu et al., 2023), which is related to the density and layout distributions of element instances in the surrounding environment. Obstruction is fundamentally a problem of the relative position between the camera and the target element instances. While SVI can be collected using relatively standardized procedures, the distance and angle between environmental elements and the camera lens can vary significantly (Zou and Wang, 2022; Lumnitz et al., 2021). In studies focused on building assessments, some retrieved SVI may only show the sides or partial views of buildings due to the variation of horizontal angles, which might not provide sufficient façade features for classification or evaluation tasks (Zou and Wang, 2022). Furthermore, suburban neighborhoods with sparse, wide roads may differ significantly in openness from densely packed commercial areas in city centers. This will affect not only the quantity and distribution of SVI collected, but also the completeness and frequency with which SVI captures specific environmental elements and instances.
In summary, the complexity of the urban environment introduces additional uncertainty in SVI coverage of environmental elements. However, compared to data quality problems or impact from SVI usage practices discussed in previous sections, it has not been given enough attention. There is a notable research gap concerning the extent to which SVI accurately covers the spatial instances in the environment, corresponding to the visual elements it aims to represent in the image space. Limited and relevant examples are only about the greenness visibility (Yan et al., 2023; Labib et al., 2021), where eye-level greenness (from SVI) and Viewshed Greenness Visibility Index (from GIS simulation) are compared. The focuses are about the similarity and discrepancies between the two indicators, rather than element coverage potential of SVI. For this reason, new perspective and method are explored to estimate the SVI coverage on urban environmental elements in this study, quantitatively incorporating the impact from environmental complexity. The details are depicted in the following sections.
@@ 3 Methodology
Figure 1: Research framework. The isovist analysis example is generated based on ‘t4gpd’ Python library.
@@ 3.1.2 Step 1 – Coverage Estimation based on Geometric Analysis
Figure 2: A simplified workflow integrating isovist analysis and computer vision technology for estimating the SVI coverage on building facades.
A Python script is designed to carry out simplified isovist analysis for SVI location points. As shown in the workflow illustrated in Figure 2, sampling points in 2 m interval are extracted from the boundaries of building footprints, as the unit representation of building facades which are potentially visible. A threshold of 50 m is set for isovist analysis, as the proximity of the maximum distance where human can achieve an efficient observation in complex urban environment. For each SVI location, lines of sight are first constructed towards all the building samplings points within the distance threshold. The spatial join method in ‘GeoPandas’ Python library is then applied to filter lines of sight which are not intersected with the surrounding building footprints. Each line of sight filtered counts as once a building sampling point can be seen via a specific SVI location. While there are existing tools to carry out 2D isovist analysis (Leduc, 2024), such as ‘t4gpd’ [^1] https://github.com/thomas-leduc/t4gpd/, they are not fully applicable in our work due to being computationally intensive for this kind of analysis. Thus, we develop our own implementation.
@@ 3.1.3 Step 2 – Coverage Validation based on Image Content
To validate whether the lines of sight are blocked by non-building elements in the real street environment, we calculate the absolute angles for lines of sight compared to the true north, and relocate them in the image space of SVI, with the SVI metadata of location and heading. For each SVI, semantic segmentation is conducted via the Python library ‘ZenSVI’ [^2] https://github.com/koito19960406/ZenSVI/, to detect different street view elements, such as building, road, sky, vehicle and greenery. Then the SVI as a panoramic image is horizontally divided into 12 pieces with equal angle range. Sinusoidal projection is applied to each angle range to restore the distortion of street view elements at polar points and suit for an eye-level view. The proportion of building elements with respect to the non-sky and non-ground elements in the angle range are calculated, as shown in Formula 5. For angle ranges with building proportion below a certain threshold, the lines of sight within the ranges can be removed, as they are highly likely to be blocked by other street obstacles. By summarizing the remaining lines of sight by building footprints and by local geospatial units, we gain a cumulative distribution of SVI coverage potential. More comprehensive indicators can be calculated to describe the SVI coverage on building facades in different dimensions.
@@ 4.1 Data Collection
The largest and most popular commercial street view service – Google Street View (GSV) is selected as the source of SVI data in this case study. We achieve a thorough search for all the latest SVI locations in the Greater London administrative area, via the official Google Maps API and the python library ‘streetlevel’ [^3] https://github.com/sk-zk/streetlevel/. ‘Streetlevel’ provides a feasible method to fetch all the available SVI locations by map tiles. The method can succeed the traditional SVI collection method relying on road sampling points, which may result in an incomplete SVI searching. Totally, 2,590,604 SVI location points and their heading directions are collected in the Greater London administrative area. We randomly sampled 1% of these location points and analyzed their nearest distances to neighboring points. It was revealed that over 71.3% of the sampled points had at least one neighboring point within a 10-meter buffer. In addition to SVI data, the research adapts the building footprint and road network data from OSM as the representation of urban environment elements. Land use data from the Colouring Cities Research Programme (CCRP) (Hudson, 2024) is applied to supplement building data from OSM, providing representations of building functional types, since such information is not always available in OSM (Biljecki et al., 2023a). Detailed classification information is available in D.
@@ 4.2 Data Management
Figure 3: The data collection and pre-processing workflow in the case study. Source of the base map data: OpenStreetMap, Greater London Authority.
@@ 4.3.3 Exploring the Impact of Collection Interval on SVI Coverage
In the Experiment 3, we aim to investigate the stability of SVI coverage on building facades in terms of different SVI collection intervals. This investigation is based on the hypothesis that higher collection interval will increase both the completeness and frequency of SVI coverage on the built environment, enhancing the information density, but may not suit best for urban research due to the extra redundancies introduced and the uncertainty in distribution.
@@
On this basis, Experiment 3 explores whether there is an optimal SVI collection interval helping improve the reliability of SVI based urban research. Non-linear functions can be fitted to precisely describe the variance of CoC-B and FoC-B indicators along different intervals, respectively. By analysing the two fitted functions, especially the speed of indicator increase or decrease relative to the SVI collection interval change, it is hypothesized that we can identify certain intervals which enables sufficient completeness of SVI coverage on built environment information while helping eliminate the unnecessary redundancy and uncertainty.
@@ 4.4 Parameters and Settings
For the standardization of the study and convenience of expression, the Experiment 1 and 2 are carried out based on SVI locations resampled at 50-meter intervals from the total SVI locations searched. 50-meter’s searching radius is adopted in the isovist analysis and 50% of the building element proportion is applied to decide whether SVI achieve a coverage on building facade on the corresponding directions, thus filtering the isovist analysis results. For the Experiment 3, SVI collection intervals increasing from 10 m to 95 m in 5 m increments are applied to resample the SVI locations. For each interval, the isovist analysis radius is set differently in 30 m, 40 m, and 50 m for comparing the experiment results. The same 50% of the building element proportion serves as the filter for isovist analysis results.
@@ 5.1.1 Coverage Completeness Distribution by Building Functions and Sizes
In Experiment 1, the study first applies CoC-B indicator to investigate the completeness of SVI coverage on individual buildings. As a preliminary result, Figure 4 illustrates the spatial distribution of CoC-B indicator estimated in the case study area. Notable heterogeneity is observed in this completeness indicator across individual buildings, which varies according to building sizes, and locations. Additionally, around 37.6% of total buildings are recorded with CoC-B equal to 0, indicating that SVI may fail to reach these buildings within the threshold distance, or the buildings are blocked visually by other street elements in SVI.
Figure 4: Map illustrating the distribution of CoC-B indicator across the case study area.
Figure 5: Boxplots showing the distribution of CoC-B for different function types and size quantiles.
@@ 5.1.2 SVI Covered Building Information and Its Original Distribution
The study first identifies dominant buildings and building types within each H3 level-9 grid, by mapping the top 10% of buildings with highest FoC-B in the grids, and coloring the grids according to the building types with highest FoC-A values, as illustrated in Figure 6. Buildings with traffic, community services, industrial and business functions, along with a multitude of unclassified, large-scale buildings, emerge as the individual buildings more frequently viewed and having a larger visual impact on the local areas. However, when analyzing dominant building types at the grid level, we find that non-residential building types are predominantly visible only within the City of London, the central area of the case study, and in a few isolated grids. Beyond these, residential buildings are the most frequently viewed and serve as general background in the building information covered by most grids in the study area. According to Table 4 and the method depicted in G, in total there are 62.2% residential buildings covered in SVI, which correspond to 66.2% residential population in the study area.
Figure 6: Mapping the dominant building instance and building types that show highest frequency to view via SVI in each local area.
@@
Figure 7: Scatter plots showing the association between FoC-A indicator of a specific building type and count proportion of the building type in H3 level-9 grids. The correlation coefficient and the formula of the fitted trend line are labeled on each subplot.

### v3

@@ Abstract
{textblock*}
(0cm,-11cm)
This is the Accepted Manuscript version of an article published by Elsevier in the journal Computers, Environment and Urban Systems in 2025, which is available at: https://doi.org/10.1016/j.compenvurbsys.2025.102253 Cite as: Fan Z, Feng CC, Biljecki F (2025): Coverage and bias of street view imagery in mapping the urban environment. Computers, Environment and Urban Systems, 117: 102253.
{textblock*}
1.5(-0.8cm,16.3cm) © 2025, Elsevier. Licensed under the Creative Commons Attribution-NonCommercial-NoDerivatives 4.0 International (http://creativecommons.org/licenses/by-nc-nd/4.0/) Street View Imagery (SVI) has emerged as a valuable data form in urban studies, enabling new ways to map and sense urban environments. However, fundamental concerns regarding the representativeness, quality, and reliability of SVI remain underexplored, e.g. to what extent can cities be captured by such data and do data gaps result in bias. This research, positioned at the intersection of spatial data quality and urban analytics, addresses these concerns by proposing a novel and effective method to estimate SVI’s element-level coverage in the urban environment. The method integrates the positional relationships between SVI and target elements, as well as the impact of physical obstructions. Expanding the domain of data quality to SVI, we introduce an indicator system that evaluates the extent of coverage, focusing on the completeness and frequency dimensions. Taking London as a case study, three experiments are conducted to identify potential biases in SVI’s ability to cover and represent urban environmental elements, using building facades as an example. It is found that despite their high availability along urban road networks, Google Street View covers only 62.4 % of buildings in the case study area. The average facade coverage per building is 12.4 %. SVI tends to over-represent non-residential buildings, thus possibly resulting in biased analyses, and its coverage of environmental elements is position-dependent. The research also highlights the variability of SVI coverage under different data acquisition practices and proposes an optimal sampling interval range of 50–60 m for SVI collection. The findings suggest that while SVI offers valuable insights, it is no panacea — its application in urban research requires careful consideration of data coverage and element-level representativeness to ensure reliable results.
@@ 2.2.1 SVI data quality: availability, image quality, and others
SVI data quality problems can be regarded as an inherent source of uncertainties for SVI in representing urban environment. Among them, the data availability have received more attention, and there are uneven availability distribution of SVI images and services across various geographic scales. Notably, a significant number of cities worldwide still lack SVI services. Cities in Europe and North America enjoy broader SVI coverage, whereas economically underdeveloped regions in Latin America and Africa experience sparse and limited distribution (Bendixen et al., 2023; Hou and Biljecki, 2022; Quinn and Alvarez León, 2019). Moreover, there is spatial heterogeneity of SVI collection within cities where SVI services are available. For example, streets in areas characterized by high traffic volumes, dense populations, or wealthier demographics are more likely to be imaged (Fry et al., 2020). Conversely, smaller towns and rural areas are often overlooked (Szczepańska and Pietrzyk, 2020). Additionally, image availability in informal urban sectors can be compromised by the absence of accessible roads (Chen et al., 2022; He and Li, 2021), which results in the structured missingness problem (Mitra et al., 2023). In their review papers, Kang et al. (2020) and Biljecki and Ito (2021) point out that SVI is predominantly collected along streets, making it challenging to analyze variations within neighborhood built environments, potentially leading to structural issues in the completeness of information.
@@ 2.2.2 Common practices in SVI utilization
Beyond the image sampling methods and intervals, researchers also consider the impact from other parameter selections and practices on SVI utilization. Specifically, these studies have compared different SVI orientations (Kim et al., 2021), different collection positions (lanes and sidewalks) (Ki et al., 2023; Ito et al., 2024a), and different image sources and forms (crowdsourced and commercial; perspective and panoramic) (Biljecki et al., 2023b), to see if they affect how SVI summarizes and reflects the same urban environment elements, such as buildings, greenery, and sky. It is reported that though reliability of a single, crowdsourced imagery is comparable to commercial panoramas (Biljecki et al., 2023b), there are significant measurement errors for sidewalks, greenery, and roads between pedestrian and vehicle views (Ki et al., 2023). Additionally, Liu and Sevtsuk (2024) discussed the issues such as lack of clear technical definitions in street attributes extracted from SVI, the lag in CV model performance, and the absence of benchmarks. Although not the focus of this paper, these issues contribute another aspect of uncertainty in the common practices of SVI utilization.
@@ 2.2.3 Complexity of real-world environment
Additionally, when mapping buildings or trees in urban environment, obstructions can also be caused by the targeted elements themselves (Yan and Huang, 2022; Raghu et al., 2023), which is related to the density and layout distributions of element instances in the surrounding environment. Obstruction is fundamentally a problem of the relative position between the camera and the target element instances. While SVI can be collected using relatively standardized procedures, the distance and angle between environmental elements and the camera lens can vary significantly (Zou and Wang, 2022; Lumnitz et al., 2021; Huang et al., 2025). In studies focused on building assessments, some retrieved SVI may only show the sides or partial views of buildings due to the variation of horizontal angles, which might not provide sufficient façade features for classification or evaluation tasks (Zou and Wang, 2022). Furthermore, suburban neighborhoods with sparse, wide roads may differ significantly in openness from densely packed commercial areas in city centers. This will affect not only the quantity and distribution of SVI collected, but also the completeness and frequency with which SVI captures specific environmental elements and instances.
In summary, the complexity of the urban environment introduces additional uncertainty in SVI coverage of environmental elements. However, compared to data quality problems or impact from SVI usage practices discussed in previous sections, it has not been given enough attention. There is a notable research gap concerning the extent to which SVI accurately covers the spatial instances in the environment, corresponding to the visual elements it aims to represent in the image space. Limited and relevant examples are only about the greenness visibility (Yan et al., 2023; Labib et al., 2021), where Green View Index (from SVI) and Viewshed Greenness Visibility Index (from GIS simulation) are compared. The focuses are about the similarity and discrepancies between the two indicators, rather than element coverage potential of SVI. For this reason, new perspective and method are explored to estimate the SVI coverage on urban environmental elements in this study, quantitatively incorporating the impact from environmental complexity. The details are depicted in the following sections.
@@ 3 Methodology
Fig. 1: Research framework. The isovist analysis example is generated based on ‘t4gpd’ Python library.
@@ 3.1.2 Step 1 – coverage estimation based on geometric analysis
Fig. 2: A simplified workflow integrating isovist analysis and computer vision technology for estimating the SVI coverage on building facades.
A Python script is designed to carry out simplified isovist analysis for SVI location points. As shown in the workflow illustrated in Figure 2, sampling points in 2 m interval are extracted from the boundaries of building footprints, as the unit representation of building facades which are potentially visible. A threshold of 50 m is set for isovist analysis, as the proximity of the maximum distance where human can achieve an efficient observation in complex urban environment. For each SVI location, lines of sight are first constructed towards all the building samplings points within the distance threshold. The spatial join method in ‘GeoPandas’ Python library is then applied to filter lines of sight which are not intersected with the surrounding building footprints. Each line of sight filtered counts as once a building sampling point can be seen via a specific SVI location. While there are existing tools to carry out 2D isovist analysis (Leduc, 2024), such as ‘t4gpd’, [^1] https://github.com/thomas-leduc/t4gpd/ they are not fully applicable in our work due to being computationally intensive for this kind of analysis. Thus, we develop our own implementation.
@@ 3.1.3 Step 2 – coverage validation based on image content
To validate whether the lines of sight are blocked by non-building elements in the real street environment, we calculate the absolute angles for lines of sight compared to the true north, and relocate them in the image space of SVI, with the SVI metadata of location and heading. For each SVI, semantic segmentation is conducted via the Python library ‘ZenSVI’, [^2] https://github.com/koito19960406/ZenSVI/ to detect different street view elements, such as building, road, sky, vehicle and greenery (Ito et al., 2024b). Then the SVI as a panoramic image is horizontally divided into 12 pieces with equal angle range. Sinusoidal projection is applied to each angle range to restore the distortion of street view elements at polar points and suit for an eye-level view. The proportion of building elements with respect to the non-sky and non-ground elements in the angle range are calculated, as shown in Formula 5. For angle ranges with building proportion below a certain threshold, the lines of sight within the ranges can be removed, as they are highly likely to be blocked by other street obstacles. By summarizing the remaining lines of sight by building footprints and by local geospatial units, we gain a cumulative distribution of SVI coverage potential. More comprehensive indicators can be calculated to describe the SVI coverage on building facades in different dimensions.
@@ 4.1 Data collection
The largest and most popular commercial street view service – Google Street View (GSV) is selected as the source of SVI data in this case study. We achieve a thorough search for all the latest SVI locations in the Greater London administrative area, via the official Google Maps API and the python library ‘streetlevel’. [^3] https://github.com/sk-zk/streetlevel/ ‘Streetlevel’ provides a feasible method to fetch all the available SVI locations by map tiles. The method can succeed the traditional SVI collection method relying on road sampling points, which may result in an incomplete SVI searching. Totally, 2,590,604 SVI location points and their heading directions are collected in the Greater London administrative area. We randomly sampled 1 % of these location points and analyzed their nearest distances to neighboring points. It was revealed that over 71.3 % of the sampled points had at least one neighboring point within a 10-m buffer. In addition to SVI data, the research adapts the building footprint and road network data from OSM as the representation of urban environment elements. Land use data from the Colouring Cities Research Programme (CCRP) (Hudson, 2024) is applied to supplement building data from OSM, providing representations of building functional types, since such information is not always available in OSM (Biljecki et al., 2023a). Detailed classification information is available in D.
@@ 4.2 Data management
Fig. 3: The data collection and pre-processing workflow in the case study. Source of the base map data: OpenStreetMap, Greater London Authority.
@@ 4.3.3 Exploring the impact of collection interval on SVI coverage
In the Experiment 3, we aim to investigate the stability of SVI coverage on building facades in terms of different SVI collection intervals. This investigation is based on the hypothesis that smaller collection interval will increase both the completeness and frequency of SVI coverage on the built environment, enhancing the information density, but may not suit best for urban research due to the extra redundancies introduced and the uncertainty in distribution.
@@
On this basis, Experiment 3 explores whether there is an optimal SVI collection interval helping improve the reliability of SVI based urban research. Non-linear functions can be fitted to precisely describe the variation of CoC-B and FoC-B indicators along different intervals, respectively. By analyzing the two fitted functions, especially the speed of indicator increase or decrease relative to the SVI collection interval change, it is hypothesized that we can identify certain intervals which enables sufficient completeness of SVI coverage on built environment information while helping eliminate the unnecessary redundancy and uncertainty.
@@ 4.4 Parameters and settings
For the standardization of the study and convenience of expression, the Experiment 1 and 2 are carried out based on SVI locations resampled at 50-m interval from the total SVI locations searched. 50-m’s searching radius is adopted in the isovist analysis and 50 % of the building element proportion is applied to decide whether SVI achieve a coverage on building facade on the corresponding directions, thus filtering the isovist analysis results. For the Experiment 3, SVI collection intervals increasing from 10 m to 95 m in 5 m increments are applied to resample the SVI locations. For each interval, the isovist analysis radius is set differently in 30 m, 40 m, and 50 m for comparing the experiment results. The same 50 % of the building element proportion serves as the filter for isovist analysis results.
@@ 5.1.1 Coverage completeness distribution by building functions and sizes
In Experiment 1, the study first applies CoC-B indicator to investigate the completeness of SVI coverage on individual buildings. As a preliminary result, Figure 4 illustrates the spatial distribution of CoC-B indicator estimated in the case study area. Notable heterogeneity is observed in this completeness indicator across individual buildings, which varies according to building sizes and locations. Additionally, around 37.6 % of total buildings are recorded with CoC-B equal to 0, indicating that SVI may fail to reach these buildings within the threshold distance, or the buildings are blocked visually by other street elements in SVI.
Fig. 4: Map illustrating the distribution of CoC-B indicator across the case study area.
Fig. 5: Boxplots showing the distribution of CoC-B for different function types and size quantiles.
@@ 5.1.2 SVI covered building information and its original distribution
The study first identifies dominant buildings and building types within each H3 level-9 grid, by mapping the top 10 % of buildings with highest FoC-B in the grids, and coloring the grids according to the building types with highest FoC-A values, as illustrated in Figure 6. Buildings with traffic, community services, industrial and business functions, along with a multitude of unclassified, large-scale buildings, emerge as the individual buildings more frequently viewed and having a larger visual impact on the local areas. However, when analyzing dominant building types at the grid level, we find that non-residential building types are predominantly visible only within the City of London, the central area of the case study, and in a few isolated grids. Beyond these, residential buildings are the most frequently viewed and serve as general background in the building information covered by most grids in the study area. According to Table 4 and the method depicted in G, in total there are 62 % residential buildings covered in SVI, which correspond to 66.2 % residential population in the study area.
Fig. 6: Mapping the dominant building instance and building types that show highest frequency to view via SVI in each local area.
@@
Fig. 7: Scatter plots showing the association between FoC-A indicator of a specific building type and count proportion of the building type in H3 level-9 grids. The correlation coefficient and the formula of the fitted trend line are labeled on each subplot.